OpenStudy (anonymous):

You attend an amusement park with your family. Your parents buy you an all-ride pass for $20, shown as f(x). Instead of getting a pass, your parents decide to pay $4.00 for each ride they take, shown as g(x). What function shows the correct combination of these two functions to represent the total cost to them of attending the amusement park that day, shown as h(x)?

OpenStudy (anonymous):

Do you have the graphs that are the answer choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(x) = 20x, g(x) = 4, h(x) = 20x + 4 f(x) = 20, g(x) = 4, h(x) = 4 + 20 f(x) = 20, g(x) = 4x, h(x) = 4x + 20 f(x) = 20x, g(x) = 4x, h(x) = 20x + 4x

OpenStudy (anonymous):

The last one is correct

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Actually once I read the problem again I believe the answer is the third one. Since x should represent the number of rides taken and f(x) is an all ride pass, assuming this means you have to pay only once and take as many rides as you want, the number of rides would not add to the total value. Whereas g(x) is 4x since they had to pay for every ride